OpenStudy (anonymous):

Pleaseeee help if you can!! using the differential equation dN/dt=kN, where N is absorbance and t is time, derive an equation relating absorbance to time: time: 14.00 absorbance:0.20 time:16.30 absorbance:0.50

OpenStudy (anonymous):

@sloths

OpenStudy (anonymous):

Are there options or no?

OpenStudy (anonymous):

@sloths its to measure absorbance of a microorganism at different time periods. the k is a constant but that is the only equation and numbers ive been given

OpenStudy (anonymous):

Ok that's fine, but is it multiple choice?

OpenStudy (anonymous):

nope no multiple choice

OpenStudy (anonymous):

im in a bit over my head here.. what is d for?

OpenStudy (anonymous):

same, i think d means change in, so change in time or change in absorbance @sloths

OpenStudy (anonymous):

ok so i'll explain it as best i tell

OpenStudy (anonymous):

please if you can!

OpenStudy (anonymous):

it looks like you have to find the difference between first and second test, like take the first time and subtract it from the second one, 16.3-14 and put that on bottom so far we have \[\frac{ dn }{ 2.3}=kn\] then you'd take the difference of the absorbance, 0.5-0.2. then you put it on top and get this: \[\frac{ 0.3 }{ 2.3 }=kn\] thats really all i understand. did that help at all??

OpenStudy (anonymous):

yes i think ive understood it finally, thank you so much!! @sloths

OpenStudy (anonymous):

sure any time! medal me maybe?

OpenStudy (anonymous):

how do you do it? I'm new here haha @sloths

OpenStudy (anonymous):

haha same i've been here two days.. you click best response on the right side of any of my answers

OpenStudy (anonymous):

\[\frac{dN}{dt}=kN\] This equation is separable: \[\frac{dN}{N}=k~dt\] Integrate both sides: \[\begin{align*}\int\frac{dN}{N}&=\int k~dt\\ \ln|N|&=kt+C\\ e^{\ln|N|}&=e^{kt+C}\\ N&=e^{kt+C}&\text{assuming absorbance is always positive}\\ N&=e^Ce^{kt}\\ N&=Ce^{kt}&\text{since }e^{C}\text{ is itself a constant} \end{align*}\] At time \(t=14\), the absorbance is \(N=0.2\); at \(t=16.3\), \(N=0.5\). So, you have two equations: \[\begin{cases}0.2=Ce^{14k}\\ 0.5=Ce^{16.3k}\end{cases}\] Solve for \(C\) and \(k\).